OpenStudy (anonymous):

What is the slope of the line x + 5y = 9?

OpenStudy (anonymous):

Slope Intercept Form: \[y = mx + b\] Where: m is the slope b is the y-intercept

OpenStudy (mertsj):

Solve the equation for y

OpenStudy (anonymous):

Start with -x on each side

OpenStudy (anonymous):

5y=x+9

OpenStudy (anonymous):

divide by 5

OpenStudy (mertsj):

If you subtract x from the right side, how can the x be positive.

OpenStudy (anonymous):

not quite.. subtract the x from both sides to isolate the y so you can get to the form of y=mx+b :)

OpenStudy (anonymous):

Someone come help me.

OpenStudy (anonymous):

thats what i said...im trying to do step by step

OpenStudy (anonymous):

what @Mertsj said. if you are subtracting, you can't get a positive x.

OpenStudy (anonymous):

i dont want to just give the answer, its better to learn

OpenStudy (anonymous):

subtract x from what someone said divide by 5? divide what by 5 and subtract x from what??

OpenStudy (mertsj):